Isabelle is 5 years older than her brother joseph ,in 15years the sum of their age is 65. Find their present age
Found 2 solutions by math_helper, Alan3354:
Answer by math_helper(2461)

Let n = Isabelle's age

(n+15) + (n-5+15) = 65

Solve for n (Isabelle's age) then n-5 is Joseph's age.

Answer by Alan3354(69443)

Isabelle is 5 years older than her brother joseph. In 15years the sum of their age is 65. Find their present age
======================
in 15 years the sum of their ages is 65 ---> sum now is 35
---
I + J = 35
I - J = 5
------------------- Add
2I = 40
I = 20
J = 15